OpenStudy (anonymous):

which description means the same as the limit expression lim f(x) = negative infinity a. the graph falls on the left side b. the graph rises on the right side c. the graph falls on the right side d. the graph rises on the left side

OpenStudy (goformit100):

Giving you an Energetic and Warm Welcome to One & Only Open Study ............. Sir/Ma'am @shelbbss

OpenStudy (anonymous):

If you have a limit, you should also have a value that the independent variable approaches. \[\lim f(x)~\text{ means nothing, however }\lim_{x\to c}f(x)\text{ does.}\] At any rate, it looks like you would have either \(x\to-\infty\) or \(x\to+\infty\), which asks you to examine the behavior of the graph to the left or right, respectively. If the limit itself is \(-\infty\), then that means the graph "falls"